OpenStudy (anonymous):

An adventure game has a number cube with 12 equal-size faces, each labeled with a number from 1 to 12. What is the probability of rolling a 4 or a multiple of 3? Express your answer in simplest form.

OpenStudy (ilikeramen):

Hi .

OpenStudy (anonymous):

im naruto i love ramen lol

OpenStudy (faiqraees):

Rolling a 4 = only outcome is 4 = 1/12 Rolling a multiple of 3 = outcomes are 3 ,6,9,12 = 4/12 = 1/3

OpenStudy (anonymous):

umm bruh i have to sum both of those answers into one

OpenStudy (mysticfury):

but dont you gotta add the outcome of 4 in there too @FaiqRaees

OpenStudy (mysticfury):

1/3 is only the outcome of the multiples of 3

OpenStudy (anonymous):

1/12 divided by 1/3

OpenStudy (faiqraees):

@MysticFury Oh yes thank you for reminding it Rolling a 4 = only outcome is 4 = 1/12 Rolling a multiple of 3 = outcomes are 3 ,6,9,12 = 4/12 = 1/3 Probability = Rolling a 4 + Rolling a multiple of 3 Probability = 1/12 +1/3

OpenStudy (anonymous):

divide or add

OpenStudy (mysticfury):

so 5/12 thats what i was thinking , but i wanted to make sure it was correct, the put it in simplest form was the thing that made me think it was wrong since 5/12 is already in simplest form

OpenStudy (anonymous):

when i added i got 1/4

OpenStudy (anonymous):

divided

OpenStudy (mysticfury):

you dont divide it , you add it

OpenStudy (anonymous):

oo soo i guess ill give it to you

OpenStudy (anonymous):

bye